I need some help PLEASE!!!!
solve each of the following by substitution
8x-4y=16
y=2x-4
Thanks
Heather

Answer by tutorcecilia(2152)


Substitute y=2x-4 back into the first equation and solve for the x-term
8x-4(2x-4)=16
8x-8x+16=16
0+16=16
=16-16
0=0
Solution: Infinate solutions; same line